Changes

17 bytes added ,  20:12, June 12, 2009
spelling
Line 6: Line 6:  
</math>.
 
</math>.
   −
For example, the probability of a tossed fair coin coming up heads is &frac12; = 0.5 = 50% because there is one favourable outcome (heads) out of 2 possible outcomes (heads and tails).
+
For example, the probability of a tossed fair coin coming up heads is &frac12; = 0.5 = 50% because there is one favorable outcome (heads) out of 2 possible outcomes (heads and tails). Another common example is standard deck of 52 playing cards: there are 13 cards in each suit: spades, hearts, clubs and diamonds. After the deck is shuffled thoroughly, the probability that the top card on the deck is a heart is  <math>13 / 52 = 1 / 4 = 0.20</math>, or 20%.
−
 
  −
With a standard deck of 52 playing cards, there are 13 cards in each suit: spades, hearts, clubs and diamonds. After the deck is shuffled thoroughly, the probability that the top card on the deck is a heart is  <math>13 / 52 = 1 / 4 = 0.25</math>, or 25%.